2 years ago, you acquired a 10-year 0% coupon, $1000 face value bond at a YTM of 12%. Today, you sold this bond at a YTM of 8%. Calculate your annualized Horizon Yield [HY]

Assuming sem-annual compounding: answer 28.7842%

With a financial calculator, how do you find this?

Bonds of RCY Corporation with a face value of $1000 sells for $960, mature in 5 years, and have a 7% coupon rate paid semiannually. Calculate the investor's RCY by assuming the following:

- Bond sold to yield at 7% and the end of the 3-year holding period.
- Reinvestment rate 6% APR during this holding period.

Answers:

annual compounding 8.5058%

semi-annual compounding 8.3322%

$226.3943 = FVA (coupons)

$1,226.3943 = Total Return

How ere these found? Use financial calculator
Answer #1

1.
Step 1: Price today
N=8
PMT=0
I/Y=8%
FV=-1000
CPT PV=540.2688

Step 2: Price 2 years ago
N=10
PMT=0
I/Y=12%
FV=-1000
CPT PV=321.97324

Step 3: Annualised yield
N=2
FV=-540.2688
PMT=0
PV=321.97324
CPT I/Y=28.7842%

2.
Step 1: Price 3 years later
N=2*2=4
PMT=-7%*1000/2=-35
FV=-1000
I/Y=7%/2=3.5%
CPT PV=1000.0000

Step 2: Future value of coupons
N=3*2=6
PMT=-7%*1000/2=-35
I/Y=6%/2=3%
PV=0
CPT FV=$226.3943

Step 3: Total return
=1000.0000+226.3943=1226.3943

Step 4: Annual compounding
N=3
PMT=0
FV=-1226.3943
PV=960
CPT I/Y=8.5058%

Step 5: Semi annual compounding
=((1+8.5058%)^(1/2)-1)*2
=8.3322%